# MapDream: Task-Driven Map Learning for Vision-Language Navigation

### v1

@@ 3.2 Map-in-the-Loop Architecture
MapDream adopts a two-component architecture consisting of a task-driven map module and a VLN policy that operate throughout both training stages. At each time step $t$, the map module receives an egocentric observation history $O_{t}$, the current frame $o_{t}$, and the instruction $I$, and autoregressively generates a multi-channel BEV representation $M_{t}$ that summarizes navigationally relevant spatial context. This BEV map serves as an explicit spatial interface between perception and control and is provided to the VLN policy together with $O_{t}$, $o_{t}$, and $I$.
The VLN policy consumes the map and predicts a sequence of $N$ future actions $a_{t:t+N-1}$, enabling multi-step planning under partial observability. Across both supervised pre-training and reinforcement fine-tuning, the two components are optimized either with separate objectives or under a unified navigation reward, while remaining connected through the BEV representation, which functions as the sole spatial representation passed between modules.
@@ 4.2.2 Training Details
In Stage 2, we apply GRPO optimization with a rollout size of 8, a clipping parameter of 0.28, and a KL coefficient of 0.0. The model is trained for 2000 steps with a learning rate of $1\times 10^{-6}$, jointly fine-tuning both the map module and the VLN policy. Reinforcement signals consist of action and format rewards, enabling credit assignment to both spatial representations and policy decisions.

### v2

# Kaihui Wang3, Bo Zhang3, Zhizhong Su3, Deying Li1, Zhaoxin Fan2,‡

@@ 3.2 Map-in-the-Loop Architecture
MapDream adopts a two-component architecture consisting of a task-driven map module and a VLN policy that operate throughout both training stages. At each time step $t$, the map module receives an egocentric observation history $O_{t}$, the current frame $o_{t}$, and the instruction II, and autoregressively generates a multi-channel BEV representation $M_{t}$ that summarizes navigationally relevant spatial context. This BEV map serves as an explicit spatial interface between perception and control and is provided to the VLN policy together with $O_{t}$, $o_{t}$, and II.
The VLN policy consumes the map and predicts a sequence of NN future actions $a_{t:t+N-1}$, enabling multi-step planning under partial observability. Across both supervised pre-training and reinforcement fine-tuning, the two components are optimized either with separate objectives or under a unified navigation reward, while remaining connected through the BEV representation, which functions as the sole spatial representation passed between modules.
@@ 4.2.2 Training Details
In Stage 2, we apply GRPO optimization with a rollout size of 8, a clipping parameter of 0.28, and a KL coefficient of $0.$ The model is trained for 2000 steps with a learning rate of $1\times 10^{-6}$, jointly fine-tuning both the map module and the VLN policy. Reinforcement signals consist of action and format rewards, enabling credit assignment to both spatial representations and policy decisions.